# Automated Knowledge Concept Annotation and Question Representation Learning for Knowledge Tracing

### v1

@@ 1 Introduction
The recent years have witnessed a surge in online learning platforms (Adedoyin & Soykan, 2023; Gros & García-Peñalvo, 2023), where students learn new knowledge concepts, which are then tested through exercises. Needless to say, personalization is crucial for effective learning: it allows that new knowledge concepts are carefully tailored to the current knowledge state of the student, which is more effective than one-size-fits-all approaches to learning (Cui & Sachan, 2023; Xu et al., 2024). However, such personalization requires that the knowledge of students is continuously assessed, which highlights the need for knowledge tracing (KT).
@@
In this work, we propose a novel framework for automated knowledge concept annotation and question representation learning, which we call KCQRL [^1] Our code is available at https://github.com/oezyurty/KCQRL .. Our KCQRL framework is flexible and can be applied on top of any existing KT model, and we later show that our KCQRL consistently improves the performance of state-of-the-art KT models by a clear margin. Importantly, our framework is carefully designed to address the two limitations 1 and 2 from above. Technically, we achieve this through the following three modules:
@@
    We develop a novel, automated KC annotation approach using large language models (LLMs) that both generates solutions to the questions and labels KCs for each solution step. Thereby, we effectively circumvent the need for manual annotation from domain experts ($\rightarrow$ limitation 1).
@@ 2 Preliminaries: Standard Formulation of Knowledge Tracing
Figure 1: Overview of standard KT formulation and its limitations.
Knowledge tracing: We consider the standard formulation of KT (e.g., Piech et al., 2015; Sonkar et al., 2020; Zhou et al., 2024), which is the performance prediction of the next exercise for a student based on time-series data (see Fig. 1). Specifically, for a student $s$, the history of $t$ exercises are modeled as $s=\{e_{i}\}_{i=1}^{t}$ with exercises $e_{i}$. Each exercise $e_{i}$ is a 3-tuple, i.e. $e_{i}=\{q_{i},\,\{c_{i,j}\}_{j=1}^{N_{q_{i}}},\,r_{i}\}$, where $q_{i}\in{\mathbb{Q}}$ is the question ID, $c_{i,j}\in{\mathbb{C}}$ is a KC ID of $q_{i}$, and $r_{i}\in\{0,1\}$ is the binary response (=incorrect/correct) of student $s$ for exercise $e_{i}$.
@@
Limitations: In practice, the above KT formulation has two key limitations: 1 It requires thousands of questions from ${\mathbb{Q}}$ to be manually annotated with each relevant KC from ${\mathbb{C}}$ among hundreds of categories Choi et al. (2020b); Wang et al. (2020); Liu et al. (2023c). Domain experts must ensure the consistency of manual annotations, a process that is not only resource-intensive but also prone to human error (Moore et al., 2024). 2 Existing KT models overlook the semantic context of both questions and knowledge concepts (KCs), which weakens their ability to model learning sequences effectively. Typically, KT models initialize random embeddings $\textrm{Emb}(q_{i})$ for each question $q_{i}\in{\mathbb{Q}}$ and/or $\textrm{Emb}(c_{i,j})$ for each KC $c_{i,j}\in{\mathbb{C}}$ (Abdelrahman et al., 2023; Zhu et al., 2023). The embeddings are then trained as part of the supervised learning objective of the KT task. As a result, the parameters $\theta$ of a KT model $F_{\theta}$ are trained simultaneously for both (a) learning the relationships between questions and KCs and (b) sequential dynamics of student learning histories.
@@ 3 Proposed KCQRL Framework
Our KCQRL framework has three main modules (see Fig. 2): (1) The first module takes the question content and annotates the question with a step-by-step solution and the corresponding knowledge concepts (KCs) (Sec. 3.1). (2) The second module learns the representations of questions by leveraging solution steps and KCs via a tailored contrastive learning objective (Sec. 3.2). (3) The third module integrates the learned question representations into existing KT models (Sec.3.3).
3.1 Knowledge Concept Annotation via LLMs (Module 1)
Our KCQRL leverages the complex reasoning abilities of an LLM $P_{\phi}(\cdot)$ and annotates the KCs of a question in a grounded manner. This is done in three steps: (i) the LLM generates the step-by-step solution to reveal the underlying techniques to solve the problem. (ii) The LLM then annotates the required KCs to solve the problem, informed by the solution steps. (iii) Finally, it further maps each solution step with its underlying KCs, which is particularly important for the second component of our framework. The details of these three steps are given below: [^2] For each step, we provide the exact prompts in Appendix H.
(i) Solution step generation: For a question $q\in{\mathbb{Q}}$, our framework generates the solution steps $s_{1},\,\ldots,\,s_{n}$ via chain-of-thought (CoT) prompting (Wei et al., 2022) of $P_{\phi}(.)$. Here, each $s_{k}$ is a coherent language sequence that serves as an intermediate step towards solving the problem $q$. To solve the problems, $s_{k}$ is sampled sequentially using a decoding algorithm; in our framework, we use temperature sampling, i. e., $s_{k}\sim P_{\phi}(s_{k}\,|\,\textrm{prompt}(q),\,s_{1},\,\ldots,\,s_{k-1})$, where $s_{k}$ is the $k$-th solution step of a question $q$ and $\textrm{prompt}(q)$ is the CoT prompt for $q$. Note that CoT prompting allows us to decompose the solutions in multiple steps, which is especially relevant for questions that require complex problem-solving such as in Maths, Computer Science, Physics, Chemistry, and Medicine.
@@ 3.1 Knowledge Concept Annotation via LLMs (Module 1)
(iii) Solution step $\rightarrow$ KC mapping: Not all KCs have been practiced at each solution step of a problem. To better understand the association between each solution step and KC, our KCQRL framework further maps each solution step to its associated KCs. Specifically, $P_{\phi}(.)$ is presented with the question content $q$, the solution steps $\{s_{k}\}_{k=1}^{n}$, and the KCs $\{c_{j}\}_{j=1}^{m}$. Then, the LLM is asked to sequentially generate the relevant pairs of solution step and KC, i. e.,
$(s_{\pi_{s}(l)},c_{\pi_{c}(l)})\sim P_{\phi}((s_{\pi_{s}(l)},c_{\pi_{c}(l)})\,|\,\textrm{prompt}(q,\{s_{k}\}_{k=1}^{n},\{c_{j}\}_{j=1}^{m}),(s_{\pi_{s}(1)},c_{\pi_{c}(1)}),\ldots,(s_{\pi_{s}(l-1)},c_{\pi_{c}(l-1)})),$
@@ 3.2 Representation Learning of Questions (Module 2)
Our KCQRL framework provides a tailored contrastive learning (CL) approach to generate embeddings for questions and solution steps that are semantically aligned with their associated KCs. [^3] For an introduction to CL, we refer to (Oord et al., 2018; Chen et al., 2020; He et al., 2020). A naïve approach would be to use a pre-trained LLM, which can then generate general-purpose embeddings. However, such general-purpose embeddings lack the domain-specific focus required in education. (We provide evidence for this in our ablation studies in Sec. 5.2.) To address this, CL allows us to explicitly teach the encoder to bring question and solution step embeddings closer to the representations of their relevant KCs. After CL training, these ‘enriched’ embeddings are aggregated and used as input for the downstream KT model to improve the modeling capabilities of the latter.
@@
To train the encoder LLM $E_{\psi}(\cdot)$ with our tailored CL objective, we first compute the embeddings of question content, solution steps, and KCs via
$z^{q}_{i}=E_{\psi}(q_{i}),\qquad z^{s}_{ik}=E_{\psi}(s_{ik}),\qquad z^{c}_{ij}=E_{\psi}(c_{ij}).$ (1)
@@
$\mathcal{L}_{q}(z^{q}_{i},z^{c}_{ij})=-\log\frac{\textrm{sim}(z^{q}_{i},z^{c}_{ij})}{\textrm{sim}(z^{q}_{i},z^{c}_{ij})+\sum_{\begin{subarray}{c}i^{\prime}\in{\mathcal{B}}\\ i^{\prime}\neq i\end{subarray}}\sum_{j^{\prime}=1}^{M_{i^{\prime}}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{I}_{\left\{\mathcal{A}(c_{ij})\neq\mathcal{A}(c_{i^{\prime}j^{\prime}})\right\}}}\,\textrm{sim}(z^{q}_{i},z^{c}_{i^{\prime}j^{\prime}})},$ (2)
@@
$\mathcal{L}_{s}(z^{s}_{ik},z^{c}_{ij})=-\log\frac{\textrm{sim}(z^{s}_{ik},z^{c}_{ij})}{\textrm{sim}(z^{s}_{ik},z^{c}_{ij})+\sum_{i^{\prime}\in{\mathcal{B}}}\sum_{j^{\prime}=1}^{M_{i^{\prime}}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{I}_{\left\{\mathcal{A}(c_{ij})\neq\mathcal{A}(c_{i^{\prime}j^{\prime}})\right\}}}\,\textrm{sim}(z^{s}_{ik},z^{c}_{i^{\prime}j^{\prime}})},$ (4)
@@ 3.3 Improving Knowledge Tracing via Learned Question Embeddings (Module 3)
The final component of our KCQRL framework proceeds in a simple yet effective manner by integrating the learned representations of each question into existing KT models. To achieve this, we simply replace the randomly initialized question embeddings of the KT model $F_{\theta}$ with our learned representations. To keep the dimensionality of the embeddings consistent across the questions with different number of solution steps, we calculate the embedding of a question via
@@ 4 Experimental Setup
Baselines: Note that our framework is highly flexible and works with any state-of-the-art KT model. We thus consider a total of 15 KT models: DKT (Piech et al., 2015), DKT+ (Yeung & Yeung, 2018), KQN (Lee & Yeung, 2019), qDKT (Sonkar et al., 2020), IEKT (Long et al., 2021), AT-DKT (Liu et al., 2023a), QIKT (Chen et al., 2023), DKVMN (Zhang et al., 2017), DeepIRT (Yeung, 2019), ATKT (Guo et al., 2021), SAKT (Pandey & Karypis, 2019), SAINT (Choi et al., 2020a), AKT (Ghosh et al., 2020), simpleKT (Liu et al., 2023b). and sparseKT (Huang et al., 2023). For all the baselines, we make the following comparison: (a) the original implementations without our framework versus (b) the KT model together with our KCQRL framework. Hence, any performance improvement must be attributed to our framework.
We leverage pykt library (Liu et al., 2022b) for the implementation. We follow prior literature (e.g., Piech et al., 2015; Sonkar et al., 2020; Zhou et al., 2024) and evaluate the performance of KT models based on the AUC.
Implementation details of KCQRL: We leverage the reasoning abilities of GPT-4o [^5] The exact version is gpt-4o-2024-05-13 at https://platform.openai.com/docs/models/gpt-4o for our LLM $P_{\phi}(\cdot)$ in Sec. 3.1. We provide the exact prompts in Appendix H. We use BERT (Devlin, 2019) as our LLM encoder $E_{\psi}(\cdot)$ for representation learning in Sec. 3.2. For the elimination of false negative pairs, we use HDBSCAN (Campello et al., 2013) as the clustering algorithm $\mathcal{A}(\cdot)$ over Sentence-BERT (Reimers & Gurevych, 2019) embeddings of KCs. For training $E_{\psi}(\cdot)$, we use Nvidia Tesla A100 with 40GB GPU memory. We follow the standard five-fold cross-validation procedure to tune the parameters of the KT models $F_{\theta}(\cdot)$ in Sec. 3.3 and report results on the test set. For the KT models, we use NVIDIA GeForce RTX 3090 with 24GB GPU memory. Further details are given in Appendix D.
@@ 5.1 Prediction Performance
Main results: Table 2 reports the performance of different KT models, where we each compare in two variants: (a) without (i. e., Default) versus (b) with our framework. We find the following: (1) Our KCQRL framework consistently boosts performance across all KT models and across all datasets. The relative improvements range between 1 % to 7 %. (2) Our framework improves the state-of-the-art (SOTA) performance on both datasets. For XES3G5M, the AUC increases from 82.24 to 83.04 (+0.80), and, for Eedi, from 75.15 to 78.96 (+3.81). (3) For KT models that have otherwise a lower performance, the performance gains from our KCQRL framework are particularly large. Interestingly, this helps even low-performing KT models to reach near-SOTA performance when provided with semantically rich inputs from our framework. Takeaway: Our framework leads to consistent performance gains and achieves SOTA performance.
Table 2: Improvement in the performance of KT models from our KCQRL framework. Shown: AUC with std. dev. across 5 folds. Improvements are shown as both absolute and relative (%) values.
@@
Default w/ KCQRL Imp. (abs.) Imp. (%) Default w/ KCQRL Imp. (abs.) Imp. (%)
DKT 78.33 $\pm$ 0.06 82.13 $\pm$ 0.02 $+$3.80 $+$4.85 73.59 $\pm$ 0.01 74.97 $\pm$ 0.03 $+$1.38 $+$1.88
DKT$+$ 78.57 $\pm$ 0.05 82.34 $\pm$ 0.04 $+$3.77 $+$4.80 73.79 $\pm$ 0.03 75.32 $\pm$ 0.04 $+$1.53 $+$2.07
KQN 77.81 $\pm$ 0.03 82.10 $\pm$ 0.06 $+$4.29 $+$5.51 73.13 $\pm$ 0.01 75.16 $\pm$ 0.04 $+$2.03 $+$2.78
qDKT 81.94 $\pm$ 0.05 82.13 $\pm$ 0.05 $+$0.19 $+$0.23 74.09 $\pm$ 0.03 74.97 $\pm$ 0.04 $+$0.88 $+$1.19
IEKT 82.24 $\pm$ 0.07 82.82 $\pm$ 0.06 $+$0.58 $+$0.71 75.12 $\pm$ 0.02 75.56 $\pm$ 0.02 $+$0.44 $+$0.59
AT-DKT 78.36 $\pm$ 0.06 82.36 $\pm$ 0.07 $+$4.00 $+$5.10 73.72 $\pm$ 0.04 75.25 $\pm$ 0.02 $+$1.53 $+$2.08
QIKT 82.07 $\pm$ 0.04 82.62 $\pm$ 0.05 $+$0.55 $+$0.67 75.15 $\pm$ 0.04 75.74 $\pm$ 0.02 $+$0.59 $+$0.79
DKVMN 77.88 $\pm$ 0.04 82.64 $\pm$ 0.02 $+$4.76 $+$6.11 72.74 $\pm$ 0.05 75.51 $\pm$ 0.02 $+$2.77 $+$3.81
DeepIRT 77.81 $\pm$ 0.06 82.56 $\pm$ 0.02 $+$4.75 $+$6.10 72.61 $\pm$ 0.02 75.18 $\pm$ 0.05 $+$2.57 $+$3.54
ATKT 79.78 $\pm$ 0.07 82.37 $\pm$ 0.04 $+$2.59 $+$3.25 72.17 $\pm$ 0.03 75.28 $\pm$ 0.04 $+$3.11 $+$4.31
SAKT 75.90 $\pm$ 0.05 81.64 $\pm$ 0.03 $+$5.74 $+$7.56 71.60 $\pm$ 0.03 74.77 $\pm$ 0.02 $+$3.17 $+$4.43
SAINT 79.65 $\pm$ 0.02 81.50 $\pm$ 0.07 $+$1.85 $+$2.32 73.96 $\pm$ 0.02 75.20 $\pm$ 0.04 $+$1.24 $+$1.68
AKT 81.67 $\pm$ 0.03 83.04 $\pm$ 0.05 $+$1.37 $+$1.68 74.27 $\pm$ 0.03 75.49 $\pm$ 0.03 $+$1.22 $+$1.64
simpleKT 81.05 $\pm$ 0.06 82.92 $\pm$ 0.04 $+$1.87 $+$2.31 73.90 $\pm$ 0.04 75.46 $\pm$ 0.02 $+$1.56 $+$2.11
sparseKT 79.65 $\pm$ 0.11 82.95 $\pm$ 0.09 $+$3.30 $+$4.14 74.98 $\pm$ 0.09 78.96 $\pm$ 0.08 $+$3.98 $+$5.31
@@
Sensitivity to the number of students: Fig. 3 shows the prediction results for each KT model under a varying number of students available for training (starting from 5 % of the actual number of students). Our KCQRL significantly improves the performance of KT models for both datasets and for all % of available training data. In general, the performance gain from our framework tends to be larger for small-student settings. For instance, the performance of DKVMN on Eedi improves by almost +6 AUC even when only 5 % of the actual number of students are available for training. This underscores the benefits of our KCQRL during the early phases of online learning platforms and for specialized platforms (e.g., in-house training for companies). The results for other KT models are given in Appendix E. Takeaway: Our framework greatly helps generalization performance, especially in low-sample size settings.
@@
Multi-step ahead prediction: We now follow Liu et al. (2022b) and focus on the task of predicting a span of student’s responses given the history of interactions. Example: given the initial 60 % of the interactions, the task is to predict the student’s performance in the next 40 % of the exercises. We distinguish two scenarios (Liu et al., 2022b): (a) accumulative or (b) non-accumulative manner, where (a) means that KT models makes predictions in a rolling manner (so that predictions for the $n$-outcome are used to prediction the $(n+1)$-th) and (b) means that the KT models predict all future values at once (but assume that the predictions are independent).
@@
Figure 4: Improvement of our KCQRL in multi-step-ahead prediction. Plots show different KT models, where we vary the portion of observed learning history and predict the rest of the entire learning journey. Green [red] area shows the improvement [decline] from using our framework.

### v2

@@ Abstract
knowledge tracing, time series prediction, knowledge concept, KC annotation, representation learning, contrastive learning, large language models
@@ 1 Introduction
The recent years have witnessed a surge in online learning platforms (Adedoyin & Soykan, 2023; Gros & García-Peñalvo, 2023), where students learn new knowledge concepts, which are then tested through exercises. Needless to say, personalization is crucial for effective learning: it allows that new knowledge concepts are carefully tailored to the current knowledge state of the student, which is more effective than one-size-fits-all approaches to learning (Cui & Sachan, 2023; Xu et al., 2024). However, such personalization requires that the knowledge of students is continuously assessed, highlighting the need for knowledge tracing (KT).
@@
Figure 1: Overview of standard KT formulation and its limitations.
In this work, we propose a novel framework for automated knowledge concept annotation and question representation learning, which we call KCQRL [^1] Code is available at https://github.com/oezyurty/KCQRL .. Our KCQRL framework is flexible and can be applied on top of any existing KT model, and we later show that our KCQRL consistently improves the performance of state-of-the-art KT models by a clear margin. Importantly, our framework is carefully designed to address the two limitations 1 and 2 from above. Technically, we achieve this through the following three modules:
@@
    We develop a novel, automated KC annotation approach using LLMs that both generates solutions to the questions and labels KCs for each solution step. Thereby, we effectively circumvent the need for manual annotation from domain experts ($\rightarrow$ limitation 1).
@@ 2 Preliminaries: Standard Formulation of Knowledge Tracing
Knowledge tracing: We consider the standard formulation of KT (e.g., Piech et al., 2015; Sonkar et al., 2020; Zhou et al., 2024), which is the performance prediction of the next exercise for a student based on time-series data (see Fig. 1). Specifically, for each student, the history of $t$ exercises are modeled as $\{e_{i}\}_{i=1}^{t}$ with exercises $e_{i}$. Each exercise $e_{i}$ is a 3-tuple, i.e. $e_{i}=\{q_{i},\,\{c_{i,j}\}_{j=1}^{N_{q_{i}}},\,r_{i}\}$, where $q_{i}\in{\mathbb{Q}}$ is the question ID, $c_{i,j}\in{\mathbb{C}}$ is a KC ID of $q_{i}$, and $r_{i}\in\{0,1\}$ is the response (incorrect/correct) of a student for exercise $e_{i}$.
@@
Limitations: In practice, the above KT formulation has two key limitations: 1 It requires thousands of questions from ${\mathbb{Q}}$ to be manually annotated with each relevant KC from ${\mathbb{C}}$ among hundreds of categories (Choi et al., 2020b; Wang et al., 2020; Liu et al., 2023c). Domain experts must ensure the consistency of manual annotations, a process that is not only resource-intensive but also prone to human error (Moore et al., 2024). 2 Existing KT models overlook the semantic context of both questions and knowledge concepts (KCs), which weakens their ability to model learning sequences effectively. Typically, KT models initialize random embeddings $\textrm{Emb}(q_{i})$ for each question $q_{i}\in{\mathbb{Q}}$ and/or $\textrm{Emb}(c_{i,j})$ for each KC $c_{i,j}\in{\mathbb{C}}$ (Abdelrahman et al., 2023; Zhu et al., 2023). The embeddings are then trained as part of the supervised learning objective of the KT task. As a result, the parameters $\theta$ of a KT model $F_{\theta}$ are trained simultaneously for both (a) learning the relationships between questions and KCs and (b) sequential dynamics of student learning histories.
@@ 3 Proposed KCQRL Framework
Our KCQRL framework has three main modules (see Fig. 2): (1) The first module takes the question content and annotates the question with a step-by-step solution and the corresponding knowledge concepts (KCs) (Sec. 3.1). (2) The second module learns the representations of questions by leveraging solution steps and KCs via a tailored contrastive learning (Sec. 3.2). (3) The third module integrates the learned question representations into existing KT models (Sec.3.3).
3.1 KC Annotation via LLMs (Module 1)
Our KCQRL leverages the complex reasoning abilities of an LLM $P_{\phi}(\cdot)$ and annotates the KCs of a question in a grounded manner. This is done in three steps: (i) the LLM generates the step-by-step solution to reveal the underlying techniques to solve the problem. (ii) The LLM then annotates the required KCs to solve the problem, informed by the solution steps. (iii) Finally, it further maps each solution step with its underlying KCs, which is particularly important for the second component of our framework. The details of these three steps are given below: [^2] For each step, we provide the exact prompts in Appendix I.
(i) Solution step generation: For a question $q\in{\mathbb{Q}}$, our framework generates the solution steps $s_{1},\,\ldots,\,s_{n}$ via chain-of-thought (CoT) prompting (Wei et al., 2022) of $P_{\phi}(.)$. Here, each $s_{k}$ is a coherent language sequence that serves as an intermediate step towards solving the problem $q$. To solve the problems, $s_{k}$ is sampled sequentially using a decoding algorithm; in our framework, we use temperature sampling, i. e., $s_{k}\sim P_{\phi}(s_{k}\,|\,\textrm{prompt}(q),\,s_{1},\,\ldots,\,s_{k-1})$, where $s_{k}$ is the $k$-th solution step of a question $q$ and $\textrm{prompt}(q)$ is the CoT prompt for $q$. Note that CoT prompting allows us to decompose the solutions in multiple steps, which is especially relevant for questions that require complex problem-solving such as in Math.
@@ 3.1 KC Annotation via LLMs (Module 1)
(iii) Solution step $\rightarrow$ KC mapping: Not all KCs have been practiced at each solution step of a problem. To better understand the association between each solution step and KC, our KCQRL framework further maps each solution step to its associated KCs. Specifically, $P_{\phi}(.)$ is presented with the question content $q$, the solution steps $\{s_{k}\}_{k=1}^{n}$, and the KCs $\{c_{j}\}_{j=1}^{m}$. Then, the LLM is asked to sequentially generate the relevant pairs of solution step and KC, i. e., $(s_{\pi_{s}(l)},c_{\pi_{c}(l)})\sim P_{\phi}\big{(}(s_{\pi_{s}(l)},c_{\pi_{c}(l)})\,\big{|}\,\textrm{prompt}(q,\{s_{k}\}_{k=1}^{n},\{c_{j}\}_{j=1}^{m}),\\ (s_{\pi_{s}(1)},c_{\pi_{c}(1)}),\ldots,(s_{\pi_{s}(l-1)},c_{\pi_{c}(l-1)})\big{)},$
@@ 3.2 Representation Learning of Questions (Module 2)
Our KCQRL framework provides a tailored contrastive learning (CL) approach to generate embeddings for questions and solution steps that are semantically aligned with their associated KCs. [^3] For an introduction to CL, we refer to (Oord et al., 2018; Chen et al., 2020; He et al., 2020). A naïve approach would be to use a pre-trained LM, which can then generate general-purpose embeddings. However, such general-purpose embeddings lack the domain-specific focus required in education. (We provide evidence for this in our ablation studies in Sec. 5.2.) To address this, CL allows us to explicitly teach the encoder to bring question and solution step embeddings closer to the representations of their relevant KCs. After CL training, these ‘enriched’ embeddings are aggregated and used as input for the downstream KT model to improve the modeling capabilities of the latter.
@@
To train the encoder LM $E_{\psi}(\cdot)$ with our tailored CL objective, we first compute the embeddings of question content, solution steps, and KCs via
$z^{q}_{i}=E_{\psi}(q_{i}),\quad z^{s}_{ik}=E_{\psi}(s_{ik}),\quad z^{c}_{ij}=E_{\psi}(c_{ij}).$ (1)
@@
$\mathcal{L}_{q}(z^{q}_{i},z^{c}_{ij})=-\log\frac{\textrm{sim}(z^{q}_{i},z^{c}_{ij})}{\textrm{sim}(z^{q}_{i},z^{c}_{ij})+\sum_{\begin{subarray}{c}i^{\prime}\in{\mathcal{B}}\\ i^{\prime}\neq i\end{subarray}}\sum_{j^{\prime}=1}^{M_{i^{\prime}}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{I}_{\left\{\mathcal{A}(c_{ij})\neq\mathcal{A}(c_{i^{\prime}j^{\prime}})\right\}}}\,\textrm{sim}(z^{q}_{i},z^{c}_{i^{\prime}j^{\prime}})},$
@@
$\mathcal{L}_{s}(z^{s}_{ik},z^{c}_{ij})=-\log\frac{\textrm{sim}(z^{s}_{ik},z^{c}_{ij})}{\textrm{sim}(z^{s}_{ik},z^{c}_{ij})+\sum_{i^{\prime}\in{\mathcal{B}}}\sum_{j^{\prime}=1}^{M_{i^{\prime}}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{I}_{\left\{\mathcal{A}(c_{ij})\neq\mathcal{A}(c_{i^{\prime}j^{\prime}})\right\}}}\,\textrm{sim}(z^{s}_{ik},z^{c}_{i^{\prime}j^{\prime}})},$
@@ 3.3 Improving Knowledge Tracing via Learned Question Embeddings (Module 3)
The final component of our KCQRL framework proceeds in a simple yet effective manner by integrating the learned representations of each question into existing KT models. To achieve this, we simply replace the randomly initialized question embeddings of the KT model $F_{\theta}$ with our learned representations. To keep the dimensionality of the embeddings consistent across questions with different numbers of solution steps, we calculate the embedding of a question as
@@ 4 Experimental Setup
Translated from Chinese.
Baselines: Note that our framework is highly flexible and works with any state-of-the-art KT model. We thus consider a total of 15 KT models: DKT (Piech et al., 2015), DKT+ (Yeung & Yeung, 2018), KQN (Lee & Yeung, 2019), qDKT (Sonkar et al., 2020), IEKT (Long et al., 2021), AT-DKT (Liu et al., 2023a), QIKT (Chen et al., 2023), DKVMN (Zhang et al., 2017), DeepIRT (Yeung, 2019), ATKT (Guo et al., 2021), SAKT (Pandey & Karypis, 2019), SAINT (Choi et al., 2020a), AKT (Ghosh et al., 2020), simpleKT (Liu et al., 2023b). and sparseKT (Huang et al., 2023). For all the baselines, we make the following comparison: (a) the original implementations without our framework versus (b) the KT model together with our KCQRL framework. Hence, any performance improvement must be attributed to our framework. We follow prior literature (e.g., Piech et al., 2015; Sonkar et al., 2020; Zhou et al., 2024) and evaluate the performance of KT models based on the AUC.
Implementation details are in Appendix D.
@@ 5.1 Prediction Performance
Main results: Table 2 reports the performance of different KT models, where we each compare in two variants: (a) without (i. e., Default) versus (b) with our framework. We find the following: (1) Our KCQRL framework consistently boosts performance across all KT models and across all datasets. The relative improvements range between 1 % to 7 %. (2) Our framework improves the state-of-the-art (SOTA) performance on both datasets. For XES3G5M, the AUC increases from 82.24 to 83.04 (+0.80), and, for Eedi, from 75.15 to 78.96 (+3.81). (3) For KT models that with lower default performance, the performance gains from our KCQRL framework are particularly large. Interestingly, this helps even low-performing KT models to reach near-SOTA performance when provided with semantically rich inputs from our framework. Takeaway: Our KCQRL leads to consistent performance gains and achieves SOTA performance.
Table 2: Improvement in the performance of KT models with our KCQRL framework. Shown are AUC with std. dev. across 5 folds and improvements as relative (%) values.
@@
Default w/ KCQRL Imp. (%) Default w/ KCQRL Imp. (%)
DKT 78.33 $\pm$ 0.06 82.13 $\pm$ 0.02 $+$4.85 73.59 $\pm$ 0.01 74.97 $\pm$ 0.03 $+$1.88
DKT$+$ 78.57 $\pm$ 0.05 82.34 $\pm$ 0.04 $+$4.80 73.79 $\pm$ 0.03 75.32 $\pm$ 0.04 $+$2.07
KQN 77.81 $\pm$ 0.03 82.10 $\pm$ 0.06 $+$5.51 73.13 $\pm$ 0.01 75.16 $\pm$ 0.04 $+$2.78
qDKT 81.94 $\pm$ 0.05 82.13 $\pm$ 0.05 $+$0.23 74.09 $\pm$ 0.03 74.97 $\pm$ 0.04 $+$1.19
IEKT 82.24 $\pm$ 0.07 82.82 $\pm$ 0.06 $+$0.71 75.12 $\pm$ 0.02 75.56 $\pm$ 0.02 $+$0.59
AT-DKT 78.36 $\pm$ 0.06 82.36 $\pm$ 0.07 $+$5.10 73.72 $\pm$ 0.04 75.25 $\pm$ 0.02 $+$2.08
QIKT 82.07 $\pm$ 0.04 82.62 $\pm$ 0.05 $+$0.67 75.15 $\pm$ 0.04 75.74 $\pm$ 0.02 $+$0.79
DKVMN 77.88 $\pm$ 0.04 82.64 $\pm$ 0.02 $+$6.11 72.74 $\pm$ 0.05 75.51 $\pm$ 0.02 $+$3.81
DeepIRT 77.81 $\pm$ 0.06 82.56 $\pm$ 0.02 $+$6.10 72.61 $\pm$ 0.02 75.18 $\pm$ 0.05 $+$3.54
ATKT 79.78 $\pm$ 0.07 82.37 $\pm$ 0.04 $+$3.25 72.17 $\pm$ 0.03 75.28 $\pm$ 0.04 $+$4.31
SAKT 75.90 $\pm$ 0.05 81.64 $\pm$ 0.03 $+$7.56 71.60 $\pm$ 0.03 74.77 $\pm$ 0.02 $+$4.43
SAINT 79.65 $\pm$ 0.02 81.50 $\pm$ 0.07 $+$2.32 73.96 $\pm$ 0.02 75.20 $\pm$ 0.04 $+$1.68
AKT 81.67 $\pm$ 0.03 83.04 $\pm$ 0.05 $+$1.68 74.27 $\pm$ 0.03 75.49 $\pm$ 0.03 $+$1.64
simpleKT 81.05 $\pm$ 0.06 82.92 $\pm$ 0.04 $+$2.31 73.90 $\pm$ 0.04 75.46 $\pm$ 0.02 $+$2.11
sparseKT 79.65 $\pm$ 0.11 82.95 $\pm$ 0.09 $+$4.14 74.98 $\pm$ 0.09 78.96 $\pm$ 0.08 $+$5.31
@@
Sensitivity to the number of students: Fig. 3 shows the prediction results for two randomly selected KT models under a varying number of students available for training (starting from 5 % of the actual number of students). Our KCQRL significantly improves the performance of KT models for both datasets and for all % of available training data. In general, the performance gain from our framework tends to be larger for small-student settings. For instance, the performance of AKT on Eedi improves by +4 AUC even when only 5 % of the actual number of students are available for training. This underscores the benefits of our KCQRL during the early phases of online learning platforms and for specialized platforms (e.g., in-house training for companies). The results for other KT models are given in Appendix E. Takeaway: Our framework greatly helps generalization performance, especially in low-sample size settings.
@@
Multi-step ahead prediction: We now follow Liu et al. (2022b) and focus on the task of predicting a span of student’s responses given the history of interactions. Example: given the initial 60 % of the interactions, the task is to predict the student’s performance in the next 40 % of the exercises. We distinguish two scenarios (Liu et al., 2022b): (a) accumulative or (b) non-accumulative manner, where (a) means that KT models make predictions in a rolling manner (so that predictions for the $n$-outcome are used to prediction the $(n+1)$-th) and (b) means that the KT models predict all future values at once (but assume that the predictions are independent).
@@
Figure 4: Improvement of our KCQRL in multi-step-ahead prediction. Plots show different KT models, where we vary the portion of observed learning history and predict the rest of the entire learning journey. Green area shows the improvement from using our framework.